Suppose Mary intends to sell two software products X & Y for the next convention & budgets the following. X Y Total Units Sold. 60 40 100 Revenues, $200 $100 per unit $12,000 $ 4,000 $16,000 Variable Costs, $120 $70 per unit 7,200 2,800 10,000 Unit Contribution Margin, $80 $ 30 per unit $ 4,800 $ 1200 $ 6,000 Fixed Costs 4,500 Operating Income $ 1,500 Required: What is the BEP (in units & in Birr 


1
Expert's answer
2022-03-21T18:17:03-0400
  1. $4500
  2. (80+30+4800+1200+600)/5 = 1342
  3. 70-30= 40USD
  4. 80+30+4800+1200+600)/5 = 1342
  5. 4500-1500 = 3000.